# Local Measurement Scheme of Gravitational Curvature using Atom Interferometers

### v1

@@ II.1 Measurement of gravitational curvature
$\phi(z)=\phi_{0}+g\,z+\Gamma(z)\quad\text{with}\quad\Gamma(z)=\sum_{n=2}\frac{\phi^{(n)}}{n!}z^{n},$ (1)
@@ II.2 Idealized gravitational fields
$\phi_{\text{Ideal}}(z)=gz+\frac{1}{2}\Gamma_{0}z^{2},\quad\text{i.e.,}\quad\Gamma(z)=\frac{1}{2}\Gamma_{0}\,z^{2},$ (4)
@@ II.3 Gravitational background of the VLBAI Hannover
Note that $f$ arises from the idealized description of the CGI and should therefore yield an especially good approximation for the actual, but unknown, scale factor for small $\Delta h$. Refer to caption Figure 3: (a) Comparison of the measured phase shift $\Delta\Phi(z_{0})$ (black), the gravity gradient $\Gamma(z_{0})$ from Fig. 2 (orange), and the estimator for the gravity gradient $\hat{\Gamma}(z_{0})$ from Eq. (11) (red dashed) for three different values of $T_{R}$ (and therefore $\Delta h$), i.e. different choices of measurement resolution. (b) Phase shift magnitude for CGIs with varying baselines $\Delta h$ and corresponding mean-square error in the estimation of the gravity gradient. $\Delta\hat{\Gamma}$ is averaged over all possible initial heights in the ROI obtainable with a baseline of $\Delta h$ and $\langle\Gamma\rangle=2.75\times 10^{3}\,\mathrm{E}$ is the magnitude of the mean gravitational gradient of the facility. Fig. 3 (a) shows the excellent agreement with the correct gravity gradient $\Gamma(z_{0})$ for small baselines $\Delta h$, i.e. high spatial resolutions. We furthermore illustrate how increasing the baseline – and thus the average phase shift – results in a higher mean square error $\Delta\hat{\Gamma}$ in the estimation of the gravity gradient in Fig. 3 (b).
@@ III Discussion
We have introduced a novel differential setup to measure gravitational curvature and have simulated its behavior in a complex gravitational field. Furthermore, we defined an estimator for the gravity gradient that aligns with the true value within a 4% margin for interferometers with baselines up to $2.5\,\mathrm{m}$. The question of determining the appropriate averaging procedure to achieve alignment between the AIF phase and the gravitational signal, as in our case with the gravitational gradient case, is inherently complex, especially for arbitrary gravitational fields. Future work on this topic will be necessary, especially if one wants to measure gravitational waves or dark matter with earthbound experiments, with more elaborate interferometer geometries. This is because gravitational perturbations and non-linearities cannot be effectively shielded in those setups and must therefore be characterized with high accuracy.
Detecting changes in the gravitational gradient also has practical use cases. In the work of Stray et al. [19], a gradiometric setup of MZIs has been used to detect a tunnel, that creates a deviation of the gravity gradient on the order of $150\,\mathrm{E}$. The detection of this tunnel required a phase resolution of $17.5\,\mathrm{mrad}$, which could have been achieved with a minimal baseline in our AIF geometry of $\Delta h\approx 2\,\mathrm{m}$, cf. Fig. 3 (b). A desired phase shift magnitude of $17.5\,\mathrm{mrad}$ in the CGI would induce a single shot estimation error of the gravity gradient (of Earth) of around $100\,\mathrm{E}$, cf. Fig. 3 (b). In comparison, the gravity gradient at the VLBAI Hannover deviates along the ROI roughly $300\,\mathrm{E}$, i.e. roughly twice as much as compared to [19].

### v3

@@ II.1 Measurement of gravitational curvature
[…]
@@ II.2 Idealized gravitational fields
$\phi_{\text{Ideal}}(z)=gz+\frac{1}{2}\Gamma_{0}z^{2},\quad\text{i.e.,}\quad\Gamma(z)=\Gamma_{0},$ (4)
@@ II.3 Gravitational background of the VLBAI Hannover
Note that $f$ arises from the idealized description of the CGI and should therefore yield an especially good approximation for the actual, but unknown, scale factor for small $\Delta h$. Refer to caption Figure 3: (a) Comparison of the measured phase shift $\Delta\Phi(z_{0})$ (black), the gravity gradient $\Gamma(z_{0})$ from Fig. 2 (orange), and the estimator for the gravity gradient $\hat{\Gamma}(z_{0})$ from Eq. (11) (red dashed) for three different values of $T_{R}$ (and therefore $\Delta h$), i.e. different choices of measurement resolution. (b) Phase shift magnitude for CGIs with varying baselines $\Delta h$ and corresponding root mean-square error in the estimation of the gravity gradient. $\Delta\hat{\Gamma}$ is averaged over all possible initial heights in the ROI obtainable with a baseline of $\Delta h$ and $\langle\Gamma\rangle=2.75\times 10^{3}\,\mathrm{E}$ is the magnitude of the mean gravitational gradient of the facility. Fig. 3 (a) shows the excellent agreement with the correct gravity gradient $\Gamma(z_{0})$ for small baselines $\Delta h$, i.e. high spatial resolutions. We furthermore illustrate how increasing the baseline – and thus the average phase shift – results in a higher root mean-square error $\Delta\hat{\Gamma}$ in the estimation of the gravity gradient in Fig. 3 (b).
@@ III Discussion
We have introduced a novel differential setup to measure gravitational curvature and have simulated its behavior in a complex gravitational field. Furthermore, we defined an estimator for the gravity gradient that aligns with the true value within a 1% margin for interferometers with baselines up to $3.5\,\mathrm{m}$. The question of determining the appropriate averaging procedure to achieve alignment between the AIF phase and the gravitational signal, as in our case with the gravitational gradient case, is inherently complex, especially for arbitrary gravitational fields. Future work on this topic will be necessary, especially if one wants to measure gravitational waves or dark matter with earthbound experiments, with more elaborate interferometer geometries. This is because gravitational perturbations and non-linearities cannot be effectively shielded in those setups and must therefore be characterized with high accuracy. For example, in the VLBAI facility in Hannover targeting among others high precision gravimetry, the gravity gradient changes by approximately $300\,\mathrm{E}$ over the whole baseline. Moreover, constructions such as underground tunnels, modify the gravitational field and can be detected by gravity gradient measurements as exemplified by Stray et al. [19] employing a conventional gradiometer using two MZIs. (There, deviations of about $150\,\mathrm{E}$ required a phase resolution of $17.5\,\mathrm{mrad}$.)